OpenStudy (anonymous):

Frank has devised a formula for his catering business that calculates the number of meatballs he needs to prepare. The formula is m=4a +2c, where c= number of children; m= number of meatballs; a= number of adults. Q: How many meat balls are required for a party of 25 adults and 6 children?

OpenStudy (anonymous):

m= 4(25) + 2(6) = 100 + 12 = 112

OpenStudy (anonymous):

Thanks!